# Aligning Transformers with Continuous Feedback via Energy Rank Alignment

### v2

@@ E.2 Chemical properties
E.3 Molecular alignment details
E.3.1 Unprompted molecular generation (RDKit oracles)
@@
E.3.2 Prompted molecular generation (RDKit oracles)
@@
E.4 Unprompted molecular generation (protein-ligand docking oracles)
@@
We compute metrics on the top-100 novel and unique molecules on 20000 sampled molecules (see Figure S12) from the aligned models. We compute the mean score and the internal diversity score (IntDiv) [19] computed according to the following
@@
Appendix F Details for protein language model experiments
We consider mutating the TrpB protein at 4 sites (positions 182, 183, 184, and 186) to all of the 20 standard amino acids and compute the EVMutation score [18] for all $16000020^{4}=$ 160000 sequences (see [44] for dataset). We used the standard tokenization scheme in ESM3 and during training and inference masked all non-sequence tasks [17]. ESM3 is a bidirectional transformer and so to generate mutant sequences, we first pass in the full sequence with the mutant sites masked and use the model to “unmask” these four sites simultaneously. For the sequence track, ESM3 also contains additional tokens representing non-standard residues and special cases, and we consider sequences “invalid” if the generated token does not correspond to one of the 20 standard amino acid sequences.
@@ Appendix F Details for protein language model experiments
Appendix G Computational resources
For all chemical alignment experiments, we trained on an in-house cluster with 8 Nvidia 4080 GPUs. For LLM experiments, we used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility. Jobs run on NERSC used at most 4 Nvidia A100 GPUs (either 40GB or 80GB depending on what was allocated).
Appendix H Societal and broader impacts

### v3

@@ C.2 Chemical properties
C.3 Molecular alignment details
C.3.1 ERA ablations and comparison to DPO
$\beta$ $\gamma$ Validity $\uparrow$ Top 100 Mean $\uparrow$ Top 100 Diversity $\downarrow$
10 0.01 72.00 0.932 0.129
10 0.1 74.23 0.933 0.131
100 0.01 74.71 0.934 0.129
100 0.1 75.14 0.935 0.131
Table S4: Descriptive statistics of 10,000 generated SMILES after aligning the molecular transformer model with ERA for QED maximization across several values of $\beta$ and $\gamma$.
DPO results
$\beta_{\text{DPO}}$ Validity (%) $\uparrow$ Top-100 mean $\uparrow$ Top-100 diversity $\downarrow$
100 86.91 0.936 0.131
10 80.92 0.943 0.134
1 82.23 0.943 0.156
0.1 54.46 0.947 0.174
ERA results ($\gamma=9$)
$\beta_{\text{ERA}}$ Validity (%) $\uparrow$ Top-100 mean $\uparrow$ Top-100 diversity $\downarrow$
0.1 84.78 0.939 0.149
1 85.73 0.936 0.145
10 79.07 0.932 0.141
100 88.83 0.941 0.145
Table S5: Comparison of DPO and ERA results across different $\beta$ values on the task of QED maximization.
To evaluate the effects of the parameters $\beta$ and $\gamma$ on the performance of ERA, we assess the algorithm’s performance on the task of QED maximization by ablating $\beta$ and $\gamma$ in Table S4. We observe that higher $\beta$ leads to greater dominance of the reward and increasing values for the Top-100 mean QED whereas higher $\gamma$ leads to higher validity due to the increased regularization against the reference policy, both consistent with the minimizer presented in 5.
To assess ERA against DPO, we evaluate each method on their performance on the task of QED maximization. We note that the DPO objective does not contain an explicit reward component and instead only maximizes the margins between preferred and dis-preferred samples. Additional care needs to be taken when comparing DPO and ERA hyperparameters as $\beta_{ERA}\neq\beta_{DPO}$ due to a difference in definition. Specifically, the minimizer of the DPO objective is:
$\pi_{DPO}(x,y)\propto\exp(-\beta_{DPO}^{-1}r(x,y)+\log\pi_{ref})$
Whereas for ERA, the minimizer is
$\pi_{ERA}(x,y)\propto\exp\left(\frac{-\beta_{ERA}}{1+\gamma}r(x,y)+\frac{\gamma}{1+\gamma}\log\pi_{ref}\right)$
The formulation of the DPO objective does not permit explicitly scaling the contribution of the regularization term against the prior while ERA permits this. This also implies that
$\beta_{ERA}=\frac{1+\gamma}{\beta_{DPO}}$
Furthemore, we take $\gamma$ to be relatively large since as $\gamma\to\infty$, $\frac{\gamma}{1+\gamma}\to 1$. In our comparisons, we choose $\gamma=9$. Table S5 contains descriptive statistics of 10,000 generated SMILES after aligning the molecular transformer model for QED maximization using either DPO or ERA across several matched values of $\beta$. It is arranged such that the corresponding rows can be compared directly against each other. We did not scan any $\beta<0.1$ since the validity in those cases was low for both methods. Due to parameter matching, we expect ERA and DPO to be similar here: validity is comparable, with ERA surpassing DPO at higher $\beta_{ERA}$, as well as the top-100 average QED. ERA consistently generates diverse molecules across values of $\beta_{ERA}$ as opposed to DPO, as evidenced by the lower similarity among the top-100 generations for $\beta_{ERA}=10$ and $\beta_{ERA}=100$. This again reflects the fact that the entropy-regulated ERA objective promotes diverse generations while effectively avoiding greedy policies as opposed to DPO, which focuses only on maximizing the margins between preferred and dispreferred sequences. This highlights a key advantage of ERA since good molecular diversity is a key consideration for chemical and drug discovery tasks.
C.3.2 Unprompted molecular generation (RDKit oracles)
@@
C.3.3 Prompted molecular generation (RDKit oracles)
@@
C.3.4 Unprompted molecular generation (protein-ligand docking oracles)
@@
We compute metrics on the top-100 novel and unique molecules on 20000 sampled molecules (see Figure S12) from the aligned models. We compute the mean score and the internal diversity score (IntDiv) Hu et al. [2023] computed according to the following
@@
C.3.5 Analysis of invalid generations
Especially when the regularization $\gamma=0$, we see that ERA reduces validity of the generated SMILES. There are some limitations imposed by the fact that most of our downstream characterizations cannot be run on invalid SMILES string because of the failure of RDKit to parse. To assess the failure modes of these invalid generations, we computed two measures of the diversity: an estimate of the Levenshtein distance and an estimate of the Shannon entropy for both valid and invalid molecules.
Table S10 demonstrates properties of the valid and invalid generations from a model aligned using ERA for QED maximization for $\beta=100.0$, $\gamma=0.1$. Notably, the Shannon entropy of generated tokens and the mean Levenshtein distance between all of the generated SMILES is higher for the invalid outputs than the valid ones. These metrics suggest that there is in fact less clustering around one specific region of chemical space in these failure cases.
For the same QED maximization run ($\beta=100.0$, $\gamma=0.1$) as above, we found a significant failure mode to be that of missing tokens to close branches (67.6% of all invalid generations). Furthermore, failure to properly close rings also occurs in a signifcant amount of invalid generations (30.5%). One potential failure mode that seems to be relatively insignificant is that of failing to predict the stop token before our hard limit of 500 tokens, which occurs in fewer than 1% of invalid generations. Beyond this analysis, visual inspection of several failure cases does not appear to reveal any particular motif that is repeated among these examples beyond the model attempting to generate highly branched molecules or molecules with several rings.
Generations Shannon Entropy Levenshtein Distance
Valid 2.2574 39.8225
Invalid 2.3177 56.1066
Table S10: Shannon entropy and Levenshtein distances for valid and invalid generations from a QED maximization molecular generator alignment experiment with $\beta=100.0$, $\gamma=0.1$.
Appendix D Details for protein language model experiments
We consider mutating the TrpB protein at 4 sites (positions 182, 183, 184, and 186) to all of the 20 standard amino acids and compute the EVMutation score Hopf et al. [2017] for all $20^{4}=160000$ sequences (see Yang et al. [2023] for dataset). We used the standard tokenization scheme in ESM3 and during training and inference masked all non-sequence tasks Hayes et al. [2024]. ESM3 is a bidirectional transformer and so to generate mutant sequences, we first pass in the full sequence with the mutant sites masked and use the model to “unmask” these four sites simultaneously. For the sequence track, ESM3 also contains additional tokens representing non-standard residues and special cases, and we consider sequences “invalid” if the generated token does not correspond to one of the 20 standard amino acid sequences.
@@ Appendix D Details for protein language model experiments
Appendix E Computational resources
For all chemical alignment experiments, we trained on an in-house cluster with 8 Nvidia 4080 GPUs. For ESM3 experiments, we used resources of the National Energy Research Scientific Computing Center (NERSC), a Department of Energy Office of Science User Facility. Jobs run on NERSC used at most 4 Nvidia A100 GPUs (either 40GB or 80GB depending on what was allocated).
Appendix F Societal and broader impacts